Determine if the given point is a solution:
6.x+y = 20 P (5,4)
7.x-y=20 P(30,10)
8.2x-y=20 P (5,5)
9.X+y=10 P (2,8)
10.2x+4y=10 P (1,2)

PLEASE answer this question asap. this is Homework.
tnks in advance.:)

Answer by Alan3354(69443)

Determine if the given point is a solution:
6.x+y = 20 P (5,4)
Sub 5 for x and 4 for y and see if it = 20
5 + 4 =? 20
It doesn't, so it's not a solution.
--------------------
Do the others the same way.
email via the Thank You note to check your work.
---------------------
7.x-y=20 P(30,10)
8.2x-y=20 P (5,5)
9.X+y=10 P (2,8)
10.2x+4y=10 P (1,2)